diffrentiate x=(1/3)(3t-6)^2and tell me the meaning of symbol^

Grade:11

2 Answers

Arun
5 years ago
It is written as
 
X= \frac{1}{3}(3t -6)^{2}
Hope now you understand the meaning of ^
 
Now please tell that respect to which variable we have to differentiate
Piyush
5 years ago
After differentiating the given equation the ans is= 2(3t-6)This `^` is used to denot the power.For eg 2^3= 8 (is. 2 to the power 3 is 8)In English it is named as carat.